OpenStudy (anonymous):

Info given: Rotation period: \[T=33m/s\] \[M=3\cdot10^{30}\] \[R=\cdot10^{4}\] A pulsar radiates energy: \[P=6\cdot10^{31}W\]. How much longer will the pulsars period be after 24h?

OpenStudy (anonymous):

\[K=\dfrac{1}{2}I\cdot\omega^2\] Where \[I=1.2\cdot10^{38} kg\cdot m^2 \]\[\omega=(\dfrac{33 \dfrac{m}{s}}{10^4m})=0.0033 \dfrac{rad}{s}\]

OpenStudy (anonymous):

\[K_{pulsar1}=\dfrac{1}{2}\cdot1.2\cdot10^{38}\cdot0.0033=6.53\cdot10^{32}J\]and \[K_{pulsar2}=\dfrac{6 \cdot 10^{21}}{24\cdot60\cdot60}=6.94\cdot10^{26}J\] The lost energy over 24h: \[K_{pulsar3}=K_{pulsar1}-K_{pulsar2}\]<=> \[K_{pulsar3}=6.53\cdot10^{32}J-6.94\cdot10^{26}J\]<=> \[K_{pulsar3}=K_{pulsar1}-K_{pulsar2}\]<=> \[K_{pulsar3}=6.53\cdot10^{32}J\]

OpenStudy (anonymous):

The new period should then be \[K=\dfrac{1}{2}⋅I⋅ω^2\]<=>\[K_{pulsar3}=\dfrac{1}{2}⋅1.2\cdot10^{38}⋅ω^2\] Where i solve for ω to get the new period: \[ω=0.0033 \dfrac{rad}{s}\]

OpenStudy (anonymous):

My problem is, that it's the same value as the one i started with ...

OpenStudy (anonymous):

is this from edx? :D

OpenStudy (anonymous):

ur omega is wrong T = 33ms.. not m PER SECOND as u have written.. its time period.. not linear speed!

OpenStudy (anonymous):

edx ?

OpenStudy (anonymous):

Oh you mean milliseconds ?

OpenStudy (anonymous):

yeha.. milliseconds.. MIT X.. online course? or you are actually from MIT ?

OpenStudy (anonymous):

No im not from MIT :) im a first year student from a danish university.

OpenStudy (anonymous):

I found the problem actually, it was a calculation mistake :)

OpenStudy (anonymous):

\[\omega=\dfrac{2\pi}{T}=\dfrac{2\pi}{33ms}\] would make more sense ye ?

OpenStudy (anonymous):

Solved it x)!